OpenStudy (anonymous):

How many two digits number are there such that on reversing the digits, the number changes by a multiple of \( 5 \)?

OpenStudy (anonymous):

9 ?

OpenStudy (anonymous):

How, give examples.

OpenStudy (anonymous):

id say 9 as well after working it out... they are 15,25,35,45,55,65,75,85.and 95

OpenStudy (anonymous):

you want a multiple of 5 right ? so I guess maybe 51, 52, 53, 54, 55, 56, 57, 58, 59

OpenStudy (king):

no he wants that the number should change by a multiple of 5

OpenStudy (anonymous):

*number changes by a multiple of 5*

OpenStudy (king):

the diff between original no. and changed no. should be a multiple of 5....rite FFM?

OpenStudy (anonymous):

I am not sure I guess it's of the form \( 10a+b = 5\times k \times 10b+a \), but then there will be only \( 1: 55 \) but I am not sure.

OpenStudy (king):

61?

OpenStudy (king):

61 when u reverse equals 16 so diff. so no. changes by 45 that is a multiple of 5

OpenStudy (anonymous):

like this many no.s will come, isn't it?

OpenStudy (king):

16 also so till now 16,61 @aadarsh:it is only fr 2 digit nos. read the question

OpenStudy (anonymous):

yeah,

OpenStudy (king):

there wont be many

OpenStudy (king):

27,72

OpenStudy (anonymous):

27,72

OpenStudy (king):

38,83

OpenStudy (king):

49,94

OpenStudy (anonymous):

Correct

OpenStudy (king):

only 8 numbers FFM

OpenStudy (anonymous):

yeah, so 4 pairs

OpenStudy (king):

yeah 4 pairs

OpenStudy (anonymous):

ow changes here means difference ?

OpenStudy (king):

yup